Grade 11Mechanics

A satellite is seen every 6 hours over the equator. It is known that it rotates opposite to the direction of rotation of earth. Then the angular velocity (in radian per hour)of satellite about the centre of earth will be a) pi/2 b) pi/3 c) pi/4 d)pi/8

Arnab Mandal
14 Years agoGrade 11

2 Answers

Swapnil Saxena
14 Years ago

Let the angular velocity of the satellite with respect to the earth center is (Omega).

Angular Velocity of Earth surface at equator is 2pi/(24)=pi/12

Then relative velocity of the sattelite and the surface at  equator is (Omega + pi/12)

Time of Sight= 6 hrs

Then (Omega + pi/12)6 =2pi

==> (Omega + pi/12)   =pi/3

==> Omega = pi/4 --(ans)

Arnab Mandal
14 Years ago

WHERE IS THE REPLY SWAPNIL SAXENA??